# Indoor Multi-View Radar Object Detection via 3D Bounding Box Diffusion

### v1

@@ Abstract
Multi-view indoor radar perception has drawn attention due to its cost-effectiveness and low privacy risks. Existing methods often rely on implicit cross-view radar feature association, such as proposal pairing in RFMask or query-to-feature cross-attention in RETR, which can lead to ambiguous feature matches and degraded detection in complex indoor scenes. To address these limitations, we propose REXO (multi-view Radar object dEtection with 3D bounding boX diffusiOn), which lifts the 2D bounding box (BBox) diffusion process of DiffusionDet into the 3D radar space. REXO utilizes these noisy 3D BBoxes to guide an explicit cross-view radar feature association, enhancing the cross-view radar-conditioned denoising process. By accounting for prior knowledge that the person is in contact with the ground, REXO reduces the number of diffusion parameters by determining them from this prior. Evaluated on two open indoor radar datasets, our approach surpasses state-of-the-art methods by a margin of +4.22+4.22 AP on the HIBER dataset and +11.02+11.02 AP on the MMVR dataset.
@@ 1 Introduction
Figure 1: (a) RFMask (Wu2023_RFMask) generates horizontal-view proposals with fixed-height vertical boxes; (b) RETR (Yataka2024_retr) implicitly links queries to cross-view features via decoder cross-attention; (c) DiffusionDet (Chen2023_diffusiondet) adapted to horizontal radar allows 2D denoising but needs extra pairing with fixed-height vertical boxes; (d) REXO (ours) performs diffusion directly in 3D radar space for simple, explicit cross-view association.
Radar perception has received increasing attention due to its robustness in low-light, adversarial weather, and hazardous conditions (e.g., smoke) (Peak2023_KRadar; Yao2024; Lu20a; Sun2020_MIMORadar_AutonomousDriging; Pandharipande2023_SensingMachineLearning; Skog2024_humandetection4dradar). Depending on the application, operational specifications, and downstream tasks, radar data can be represented in different forms such as sparse detection points (Mingmin2017_RFSleep; Sengupta2020_mmPose; Yang2023_mmFi), reflection heatmaps (Fadel2015_RFCapture; Zhao2018_RFPose; Wu2023_RFMask), Doppler signatures, and raw analog-to-digital converter (ADC) data, each with unique characteristics and feature granularity.
For indoor radar perception, single‑view and multi‑view heatmaps that combine horizontal (depth‑horizontal) and vertical (depth‑vertical) projections enable object detection, pose estimation, and segmentation on a 2D image plane (Zhao2018_RFPose; Lee2023_HuPR; Wu2023_RFMask; Perry2024_MMVR). RF‑Pose (Zhao2018_RFPose; Zhao2018_RFPose3D) first fused both views with a convolutional autoencoder to regress 2D human keypoints. RFMask (Wu2023_RFMask) grafted Faster R‑CNN (Ren2017_FasterRCNN) onto horizontal heatmaps: its region‑proposal network produces horizontal candidates that are paired with fixed‑height vertical windows to avoid exhaustive cross‑view association (Fig. 1 (a)). More recently, the radar detection transformer (RETR) (Yataka2024_retr) adopted the DETR (Carion2020_detr). Decoder queries simultaneously attend to both views through cross‑attention (Fig. 1 (b)) and are directly regressed to 3D bounding boxes (BBoxes) that are classified as person or background.
On the other hand, image-based object detection has been redefined as a generative denoising process, where random noisy 2D BBoxes are iteratively refined through a diffusion denoising process to yield final clean BBox predictions (Chen2023_diffusiondet). Referred to as DiffusionDet, it decouples training and inference, and generally surpasses query‑based detectors. When ported to horizontal radar heatmaps (Fig. 1 (c)), it denoises 2D boxes but still requires the fixed‑height vertical pairing used by RFMask.
@@ 2 Related Work
Learning-based methods have advanced radar detection over traditional model-based approaches (Kay98), benefiting from open large-scale radar point cloud datasets like nuScenes (caesar2020nuscenes), Oxford RobotCar (Oxford20), and RADIATE (Sheeny2021_RADIATE). Image-based and point/voxel-based backbones (He2016_ResNet; PillarNet) extract semantic features from radar detection points, generate region proposals, and localize objects. High-resolution heatmaps (e.g., K-Radar (Peak2023_KRadar), HIBER (Wu2023_RFMask), MMVR (Perry2024_MMVR)) and raw ADC data (ADCNet) have also been leveraged by previously mentioned RF-Pose (Zhao2018_RFPose), RFMask (Wu2023_RFMask), and RETR (Yataka2024_retr). CubeLearn (CubeLearn) replaces Fourier transforms with learnable modules for an end-to-end radar pipeline, while RAMP-CNN (RAMPCNN) enhances range-angle feature extraction via Doppler cues. More recently, diffusion models have been explored for radar applications (Zhang2024_DiffusionRadar; Luan2024_DiffPCSuperResol; chi2024_rfdiffusionradiosignalgeneration; Fan2024_RFVision; Wu2024_DiffRadar). Most efforts, e.g., Radar-Diffusion (Zhang2024_DiffusionRadar; Luan2024_DiffPCSuperResol) and DiffRadar (Wu2024_DiffRadar), focus on reconstructing LiDAR-like point clouds from low-resolution radar data, while mmDiff (Fan2024_RFVision) estimates and refines pose keypoints from sparse radar points via diffusion process.
@@
Diffusion models (Song2021_DDIM; Song2019_ScoreBasedModel; Rombach2022_StableDiff; Song2023_ConsistencyModel) have shown impressive results in tasks such as image and video generation (Ho2022_VideoGen; Blattmann2023_AlignYourLatent) and multi-view synthesis (Chen2023_3DView; Yu2023_ICCV_LongTerm3DView). For perception tasks, DiffusionDet (Chen2023_diffusiondet) first reformulates object detection as a generative denoising process and proposes to model the 2D BBoxes as random parameters in the diffusion process. Diffusion-SS3D (Ho2023_DiffSS3D) proposes a diffusion-based detector to enhance the quality of pseudo-labels in semi-supervised 3D object detection by integrating it into a teacher-student framework. CLIFF (Li2024_cliff) further leverages language models to enhance diffusion-based models for open-vocabulary object detection. Diffusion models are also considered for 3D object detection (XU2024_3DiffTection) in the context of LiDAR-Camera fusion (Xiang20233_DifFusionDetDM) and other tasks such as pose estimation (Tan2024_DiffRegPose) and semantic segmentation (Gu2024_DiffusionInst; Amit2022_segdiffimagesegmentationdiffusion).
@@ 3 Preliminary
such as the denoising diffusion probabilistic model (DDPM) (Ho2020_DDPM) and the denoising diffusion implicit model (DDIM) (Song2021_DDIM), define Markovian or non-Markovian forward processes by gradually adding noise to samples $\boldsymbol{x}_{0}$, e.g., image pixels,
@@
During training, a noise prediction network is trained to estimate $\boldsymbol{\epsilon}$ from $\boldsymbol{x}_{t}$ and time index $t$ by minimizing $\min_{\theta}\|\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t)-\boldsymbol{\epsilon}\|^{2}$, where $\theta$ represents the trainable weights. During inference, a random $\boldsymbol{x}_{T}$ is drawn from the standard Gaussian distribution and iteratively denoised using the trained $\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t)$ in reverse time: $\boldsymbol{x}_{T}\rightarrow\cdots\rightarrow\boldsymbol{x}_{t}\rightarrow\boldsymbol{x}_{t-1}\rightarrow\cdots\rightarrow\boldsymbol{x}_{0}$. Sampling strategies such as DDPM (Ho2020_DDPM)
@@
with $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}\left(\boldsymbol{0},\boldsymbol{I}\right)$, and DDIM (Song2021_DDIM) can be used to trade off between quality and speed. More details are in Appendix A.
@@ 4 REXO: BBox Diffusion in 3D Radar Space
DiffusionDet (Chen2023_diffusiondet) reformulates object detection as a denoising diffusion process, treating $\boldsymbol{x}_{t}$ as 2D BBox parameters instead of image pixels. As shown in Fig. 3, we extend this to multi-view radar by lifting $\boldsymbol{x}_{t}$ to 3D BBoxes in radar coordinates: $\boldsymbol{x}_{t}=\left\{{c^{t}_{x},c^{t}_{y},c^{t}_{z},w^{t},h^{t},d^{t}}\right\}^{\top}\in\mathbb{R}^{6}$, where $\left({c^{t}_{x},c^{t}_{y},c^{t}_{z}}\right)$ defines the center and $\left({w^{t},h^{t},d^{t}}\right)$ the size at time $t$ in the Cartesian $\{\mathtt{horizontal,vertical,depth}\}$ space. Conditioned on radar heatmaps $\{\boldsymbol{Y}_{\mathtt{hor}},\boldsymbol{Y}_{\mathtt{ver}}\}$, REXO performs 3D BBox diffusion in two phases (Fig. 3): 1) a forward process that adds noise to ground-truth (GT) BBoxes $\boldsymbol{x}_{0}$ to produce random $\boldsymbol{x}_{T}$ during training, and 2) a reverse process that denoises random $\boldsymbol{x}_{T}$ to estimate noise-free $\hat{\boldsymbol{x}}_{0}$ during inference. The denoised BBoxes are also projected to the 2D image plane for supervision in both radar and image domains. We describe REXO in four parts: training, inference, ground-level constraint and loss.
@@ 4.1 Training
Taking the two radar heatmaps $\boldsymbol{Y}_{\mathtt{hor}}$ and $\boldsymbol{Y}_{\mathtt{ver}}$ as inputs, a backbone network (e.g., ResNet (He2016_ResNet)) generates horizontal-view and vertical-view radar feature maps separately: $\boldsymbol{Z}_{\mathtt{hor}}=\mathtt{backbone}\left({\boldsymbol{Y}_{\mathtt{hor}}}\right)$ and $\boldsymbol{Z}_{\mathtt{ver}}=\mathtt{backbone}\left({\boldsymbol{Y}_{\mathtt{ver}}}\right)$, where learnable parameters in the $\mathtt{backbone}$ are shared across both views. Each feature map is generated as $L$ multi-scale feature maps in $\mathbb{R}^{C\times\frac{W}{sl}\times\frac{D}{sl}}$ or $\mathbb{R}^{C\times\frac{H}{sl}\times\frac{D}{sl}}$ by using feature pyramid network (Lin2017_fpn) where $C$, $s$ and $l\in\left\{{1,\cdots,L}\right\}$ represent the number of channels, downsampling ratio over the spatial dimension and the pyramid level, respectively.
@@
via a standard ROIAlign operation (He2017_maskrcnn), where $r$ denotes a fixed spatial resolution, e.g., $r=7$. At time $t$, this process yields $N_{\mathtt{train}}$ pairs of associated radar features
@@
where $\boldsymbol{f}=\mathtt{Predictor}\left({\mathtt{e}_{t},\boldsymbol{Z}_{\mathtt{radar}}^{\mathtt{crop}}}\right)$ is the time-dependent feature. $\mathtt{e}_{t}$ denotes the timestep embedding (Ho2020_DDPM) and $\mathtt{Predictor}$ denotes the time-dependent predictor (Chen2023_diffusiondet) with the radar feature and the embedding (see Appendix B for more details). Applying these offsets produces the final image‑plane box $\widehat{\boldsymbol{b}}_{\mathtt{image}}$, achieving tighter alignment without sacrificing geometric consistency.
@@ 4.4 Loss Function
To ensure consistency between the radar and image plane representations, we adopt a simplified scheme of the Tri-plane loss (Yataka2024_retr) that directly calculates the loss of 3D BBox. REXO employs the Hungarian match cost (kuhn1955_hungarian) with a geometry-aware loss function $\mathcal{L}_{\mathtt{box}}^{\mathtt{GA}}$ computed in both the 3D and 2D spaces:
@@ 5 Experiments
We demonstrate the effectiveness of REXO through evaluations on two open radar datasets: HIBER  (Wu2023_RFMask) and MMVR (Perry2024_MMVR).
@@ 5.1 Setup
We consider RFMask (Wu2023_RFMask), DETR (Carion2020_detr) and RETR (Yataka2024_retr) as baseline methods. Additionally, we evaluate a 3D extension of RFMask, referred to as RFMask3D (see Appendix D for details), that takes the two radar views as inputs for BBox prediction. Since RFMask and DETR originally compute the BBox loss only in the 2D horizontal radar plane and the 2D image plane, respectively, we follow the implementation of RETR and enhance both methods with a unified bi-plane BBox loss. Furthermore, we introduce a DETR variant with a top-$K$ feature selection, allowing it to take features from both horizontal and vertical heatmaps as input. For RETR, we set the number of object queries to $10$. To ensure a fair comparison, we also set $N_{\mathtt{train}}=10$ for REXO during training. All methods are evaluated using $M=4$ consecutive radar frames. Additional hyperparameter settings are provided in Appendix E.

### v2

# REXO: Indoor Multi-View Radar Object Detection via 3D Bounding Box Diffusion

@@ Abstract
Multi-view indoor radar perception has drawn attention due to its cost-effectiveness and low privacy risks. Existing methods often rely on implicit cross-view radar feature association, such as proposal pairing in RFMask or query-to-feature cross-attention in RETR, which can lead to ambiguous feature matches and degraded detection in complex indoor scenes. To address these limitations, we propose REXO (multi-view Radar object dEtection with 3D bounding boX diffusiOn), which lifts the 2D bounding box (BBox) diffusion process of DiffusionDet into the 3D radar space. REXO utilizes these noisy 3D BBoxes to guide an explicit cross-view radar feature association, enhancing the cross-view radar-conditioned denoising process. By accounting for prior knowledge that the person is in contact with the ground, REXO reduces the number of diffusion parameters by determining them from this prior. Evaluated on two open indoor radar datasets, our approach surpasses state-of-the-art methods by a margin of +4.22+4.22 AP on the HIBER dataset and +11.02+11.02 AP on the MMVR dataset. The REXO implementation is available at https://github.com/merlresearch/radar-bbox-diffusion.
@@ 1 Introduction
Figure 1: (a) RFMask (Wu et al. 2023) generates horizontal-view proposals with fixed-height vertical boxes; (b) RETR (Yataka et al. 2024) implicitly links queries to cross-view features via decoder cross-attention; (c) DiffusionDet (Chen et al. 2023b) adapted to horizontal radar allows 2D denoising but needs extra pairing with fixed-height vertical boxes; (d) REXO (ours) performs diffusion directly in 3D radar space for simple, explicit cross-view association.
Radar perception has received increasing attention due to its robustness in low-light, adversarial weather, and hazardous conditions (e.g., smoke) (Paek and others 2022; Yao et al. 2024; Lu et al. 2020; Sun et al. 2020; Pandharipande et al. 2023; Skog et al. 2024). Depending on the application, operational specifications, and downstream tasks, radar data can be represented in different forms such as sparse detection points (Zhao et al. 2017; Sengupta et al. 2020; Yang et al. 2023b), reflection heatmaps (Adib et al. 2015; Zhao et al. 2018a; Wu et al. 2023), Doppler signatures, and raw analog-to-digital converter (ADC) data, each with unique characteristics and feature granularity.
For indoor radar perception, single‑view and multi‑view heatmaps that combine horizontal (depth‑horizontal) and vertical (depth‑vertical) projections enable object detection, pose estimation, and segmentation on a 2D image plane (Zhao et al. 2018a; Lee et al. 2023; Wu et al. 2023; Rahman et al. 2024). RF‑Pose (Zhao et al. 2018a, b) first fused both views with a convolutional autoencoder to regress 2D human keypoints. RFMask (Wu et al. 2023) grafted Faster R‑CNN (Ren et al. 2017) onto horizontal heatmaps: its region‑proposal network produces horizontal candidates that are paired with fixed‑height vertical windows to avoid exhaustive cross‑view association (Fig. 1 (a)). More recently, the radar detection transformer (RETR) (Yataka et al. 2024) adopted the DETR (Carion et al. 2020). Decoder queries simultaneously attend to both views through cross‑attention (Fig. 1 (b)) and are directly regressed to 3D bounding boxes (BBoxes) that are classified as person or background.
On the other hand, image-based object detection has been redefined as a generative denoising process, where random noisy 2D BBoxes are iteratively refined through a diffusion denoising process to yield final clean BBox predictions (Chen et al. 2023b). Referred to as DiffusionDet, it decouples training and inference, and generally surpasses query‑based detectors. When ported to horizontal radar heatmaps (Fig. 1 (c)), it denoises 2D boxes but still requires the fixed‑height vertical pairing used by RFMask.
@@ 2 Related Work
Learning-based methods have advanced radar detection over traditional model-based approaches (Kay 1998), benefiting from open large-scale radar point cloud datasets like nuScenes (Caesar et al. 2020), Oxford RobotCar (Barnes et al. 2020), and RADIATE (Sheeny et al. 2021). Image-based and point/voxel-based backbones (He et al. 2016; Shi et al. 2022) extract semantic features from radar detection points, generate region proposals, and localize objects. High-resolution heatmaps (e.g., K-Radar (Paek and others 2022), HIBER (Wu et al. 2023), MMVR (Rahman et al. 2024)) and raw ADC data (Yang et al. 2023a) have also been leveraged by previously mentioned RF-Pose (Zhao et al. 2018a), RFMask (Wu et al. 2023), and RETR (Yataka et al. 2024). CubeLearn (Zhao et al. 2023) replaces Fourier transforms with learnable modules for an end-to-end radar pipeline, while RAMP-CNN (Gao et al. 2021) enhances range-angle feature extraction via Doppler cues. More recently, diffusion models have been explored for radar applications (Zhang et al. 2024; Luan et al. 2024; Chi et al. 2024; Fan et al. 2024; Wu et al. 2024). Most efforts, e.g., Radar-Diffusion (Zhang et al. 2024; Luan et al. 2024) and DiffRadar (Wu et al. 2024), focus on reconstructing LiDAR-like point clouds from low-resolution radar data, while mmDiff (Fan et al. 2024) estimates and refines pose keypoints from sparse radar points via diffusion process.
@@
Diffusion models (Song et al. 2021; Song and Ermon 2019; Rombach et al. 2022; Song et al. 2023) have shown impressive results in tasks such as image and video generation (Ho et al. 2022; Blattmann et al. 2023) and multi-view synthesis (Chen et al. 2023a; Yu et al. 2023). For perception tasks, DiffusionDet (Chen et al. 2023b) first reformulates object detection as a generative denoising process and proposes to model the 2D BBoxes as random parameters in the diffusion process. Diffusion-SS3D (Ho et al. 2023) proposes a diffusion-based detector to enhance the quality of pseudo-labels in semi-supervised 3D object detection by integrating it into a teacher-student framework. CLIFF (Li et al. 2024) further leverages language models to enhance diffusion-based models for open-vocabulary object detection. Diffusion models are also considered for 3D object detection (Xu et al. 2024) in the context of LiDAR-Camera fusion (Xiang et al. 2023) and other tasks such as pose estimation (Tan et al. 2024) and semantic segmentation (Gu et al. 2024; Amit et al. 2022).
@@ 3 Preliminary
such as the denoising diffusion probabilistic model (DDPM) (Ho et al. 2020) and the denoising diffusion implicit model (DDIM) (Song et al. 2021), define Markovian or non-Markovian forward processes by gradually adding noise to samples $\boldsymbol{x}_{0}$, e.g., image pixels,
@@
During training, a noise prediction network is trained to estimate $\boldsymbol{\epsilon}$ from $\boldsymbol{x}_{t}$ and time index $t$ by minimizing $\min_{\theta}\|\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t)-\boldsymbol{\epsilon}\|^{2}$, where $\theta$ represents the trainable weights. During inference, a random $\boldsymbol{x}_{T}$ is drawn from the standard Gaussian distribution and iteratively denoised using the trained $\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t)$ in reverse time: $\boldsymbol{x}_{T}\rightarrow\cdots\rightarrow\boldsymbol{x}_{t}\rightarrow\boldsymbol{x}_{t-1}\rightarrow\cdots\rightarrow\boldsymbol{x}_{0}$. Sampling strategies such as DDPM (Ho et al. 2020)
@@
with $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}\left(\boldsymbol{0},\boldsymbol{I}\right)$, and DDIM (Song et al. 2021) can be used to trade off between quality and speed. More details are in Appendix A.
@@ 4 REXO: BBox Diffusion in 3D Radar Space
DiffusionDet (Chen et al. 2023b) reformulates object detection as a denoising diffusion process, treating $\boldsymbol{x}_{t}$ as 2D BBox parameters instead of image pixels. As shown in Fig. 3, we extend this to multi-view radar by lifting $\boldsymbol{x}_{t}$ to 3D BBoxes in radar coordinates: $\boldsymbol{x}_{t}=\left\{{c^{t}_{x},c^{t}_{y},c^{t}_{z},w^{t},h^{t},d^{t}}\right\}^{\top}\in\mathbb{R}^{6}$, where $\left({c^{t}_{x},c^{t}_{y},c^{t}_{z}}\right)$ defines the center and $\left({w^{t},h^{t},d^{t}}\right)$ the size at time $t$ in the Cartesian $\{\mathtt{horizontal,vertical,depth}\}$ space. Conditioned on radar heatmaps $\{\boldsymbol{Y}_{\mathtt{hor}},\boldsymbol{Y}_{\mathtt{ver}}\}$, REXO performs 3D BBox diffusion in two phases (Fig. 3): 1) a forward process that adds noise to ground-truth (GT) BBoxes $\boldsymbol{x}_{0}$ to produce random $\boldsymbol{x}_{T}$ during training, and 2) a reverse process that denoises random $\boldsymbol{x}_{T}$ to estimate noise-free $\hat{\boldsymbol{x}}_{0}$ during inference. The denoised BBoxes are also projected to the 2D image plane for supervision in both radar and image domains. We describe REXO in four parts: training, inference, ground-level constraint and loss.
@@ 4.1 Training
Taking the two radar heatmaps $\boldsymbol{Y}_{\mathtt{hor}}$ and $\boldsymbol{Y}_{\mathtt{ver}}$ as inputs, a backbone network (e.g., ResNet (He et al. 2016)) generates horizontal-view and vertical-view radar feature maps separately: $\boldsymbol{Z}_{\mathtt{hor}}=\mathtt{backbone}\left({\boldsymbol{Y}_{\mathtt{hor}}}\right)$ and $\boldsymbol{Z}_{\mathtt{ver}}=\mathtt{backbone}\left({\boldsymbol{Y}_{\mathtt{ver}}}\right)$, where learnable parameters in the $\mathtt{backbone}$ are shared across both views. Each feature map is generated as $L$ multi-scale feature maps in $\mathbb{R}^{C\times\frac{W}{sl}\times\frac{D}{sl}}$ or $\mathbb{R}^{C\times\frac{H}{sl}\times\frac{D}{sl}}$ by using feature pyramid network (Lin et al. 2017) where $C$, $s$ and $l\in\left\{{1,\cdots,L}\right\}$ represent the number of channels, downsampling ratio over the spatial dimension and the pyramid level, respectively.
@@
via a standard ROIAlign operation (He et al. 2017), where $r$ denotes a fixed spatial resolution, e.g., $r=7$. At time $t$, this process yields $N_{\mathtt{train}}$ pairs of associated radar features
@@
where $\boldsymbol{f}=\mathtt{Predictor}\left({\mathtt{e}_{t},\boldsymbol{Z}_{\mathtt{radar}}^{\mathtt{crop}}}\right)$ is the time-dependent feature. $\mathtt{e}_{t}$ denotes the timestep embedding (Ho et al. 2020) and $\mathtt{Predictor}$ denotes the time-dependent predictor (Chen et al. 2023b) with the radar feature and the embedding (see Appendix B for more details). Applying these offsets produces the final image‑plane box $\widehat{\boldsymbol{b}}_{\mathtt{image}}$, achieving tighter alignment without sacrificing geometric consistency.
@@ 4.4 Loss Function
To ensure consistency between the radar and image plane representations, we adopt a simplified scheme of the Tri-plane loss (Yataka et al. 2024) that directly calculates the loss of 3D BBox. REXO employs the Hungarian match cost (Kuhn 1955) with a geometry-aware loss function $\mathcal{L}_{\mathtt{box}}^{\mathtt{GA}}$ computed in both the 3D and 2D spaces:
@@ 5 Experiments
We demonstrate the effectiveness of REXO through evaluations on two open radar datasets: HIBER  (Wu et al. 2023) and MMVR (Rahman et al. 2024).
@@ 5.1 Setup
We consider RFMask (Wu et al. 2023), DETR (Carion et al. 2020) and RETR (Yataka et al. 2024) as baseline methods. Additionally, we evaluate a 3D extension of RFMask, referred to as RFMask3D (see Appendix D for details), that takes the two radar views as inputs for BBox prediction. Since RFMask and DETR originally compute the BBox loss only in the 2D horizontal radar plane and the 2D image plane, respectively, we follow the implementation of RETR and enhance both methods with a unified bi-plane BBox loss. Furthermore, we introduce a DETR variant with a top-$K$ feature selection, allowing it to take features from both horizontal and vertical heatmaps as input. For RETR, we set the number of object queries to $10$. To ensure a fair comparison, we also set $N_{\mathtt{train}}=10$ for REXO during training. All methods are evaluated using $M=4$ consecutive radar frames. Additional hyperparameter settings are provided in Appendix E.